# Finite-Sample Non-Parametric Bounds with an Application to the Causal Effect of Workforce Gender Diversity on Firm Performance

### v1

@@ Abstract
Classical Manski bounds identify average treatment effects under minimal assumptions but, in finite samples, assume that latent conditional expectations are bounded by the sample’s own extrema or that the population extrema are known a priori—often untrue in firm-level data with heavy tails. We develop a finite-sample, concentration-driven band (concATE) that replaces that assumption with a Dvoretzky–Kiefer–Wolfowitz tail bound, combines it with delta-method variance, and allocates size via Bonferroni. The band extends to a group-sequential design that controls the family-wise error when the first “significant” diversity threshold is data-chosen. Applied to 945 listed firms (2015 Q2–2022 Q1) concATE shows that senior-level gender diversity raises Tobin’s $Q$ once representation exceeds $\approx 30\%$ in growth sectors and $\approx 65\%$ in cyclical sectors.
@@ 1 Introduction
Estimating causal effects in settings with partially unobserved counterfactuals is a fundamental challenge in econometrics. Whenever only one of two potential outcomes is observed for each unit, the average treatment effect (ATE) cannot be point-identified without additional assumptions. Recognizing this, a stream of research following Manski’s seminal work has developed nonparametric bounds for causal effects under minimal assumptions (Manski,, 1990, 2003). The classical Manski bounds make virtually no assumptions beyond knowing the treatment status and an outcome bound, instead asking: how large or small could the true ATE be, given the data we actually observe? While this worst-case approach guarantees partial identification under arbitrary heterogeneity and certain forms of selection on unobservables, it comes at the cost of wide intervals. In policy settings where the cost of acting on a wrong sign is high, such honesty can be preferable to a potentially misleading precise estimate. However, Manski’s bounds have a critical limitation in finite samples: they implicitly assume the unseen counterfactual outcomes lie within known extremes (e.g. the sample minima and maxima or exogenously given bounds). In practice—especially with heavy-tailed outcomes like firm performance—this assumption is often violated. If the true outcome distribution extends beyond the observed range, the traditional bounds can severely undercover the true effect or even become uninformative (infinite) when no credible global bound is available. In short, classical bounds that are valid asymptotically may fail to cover the true ATE in finite samples when outcomes are unbounded. This exposes a methodological gap: how can we conduct inference on partially identified effects without assuming away heavy-tail risks or sacrificing finite-sample validity?
@@
We address this gap by proposing a finite-sample, concentration-driven confidence band for the ATE—henceforth concATE. The concATE methodology replaces Manski’s reliance on known outcome bounds with a probabilistic concentration bound that accounts for sampling uncertainty in extreme values. Specifically, we exploit the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality (Dvoretzky et al.,, 1956), which provides a tight, distribution-free bound on the maximum discrepancy between the empirical distribution and the true cumulative distribution. By using the DKW inequality, we can guarantee with high finite-sample probability that the empirical range (or other tail statistics) bounds the latent outcome distribution. In essence, instead of assuming the sample extremes equal the population extremes, we allow a margin such that each unobserved tail probability is covered with finite-sample confidence. We then incorporate this “DKW padding” into the estimation of Manski’s upper and lower bound components. To account for sampling variability in the observable parts (like the treated and untreated outcome means), we employ standard delta-method approximations. Finally, we combine these elements using Bonferroni’s inequality to construct a simultaneous confidence band for the ATE bounds. This concATE band controls the familywise error rate for the entire interval estimate, ensuring that with (for example) 95% confidence the true ATE lies within the band.
@@
Notably, the concATE procedure remains valid under quite general conditions: we require no parametric outcome distribution, only mild tail assumptions and allow for either independent or weakly dependent observations (such as time-series panels) with appropriate mixing conditions. The resulting inference is robust in finite samples, avoiding the need for large-sample approximations or unknown nuisance constants that plague fully nonparametric approaches. By construction, concATE eliminates the strong functional form and ignorability assumptions that conventional regression-based analyses demand, delivering credible inference even when treatment assignment may be endogenous or outcomes are highly non-normal. In contrast to an OLS or panel regression that produces a single point estimate under strict assumptions (Angrist and Pischke,, 2009), our approach yields a range of plausible effects consistent with the data and lets the data speak when identification is weak. In summary, concATE provides a new tool for causal inference under partial observability, offering the transparency of Manski’s bounds with greater practical applicability in finite samples.
@@
To demonstrate the utility of our approach, we apply it to the question: Does greater senior-level gender workforce diversity causally improve firm performance? This question has taken on renewed importance as many firms have invested heavily in Diversity, Equity, and Inclusion (DEI) initiatives, yet establishing causality is difficult due to non-random adoption of diversity practices. A rich literature in management and economics has examined links between top management team composition and organizational outcomes. The foundational “upper echelons” theory of Hambrick and Mason, (1984) posits that a firm’s strategies and performance reflect the backgrounds of its senior executives. Consistent with this view, numerous studies document associations between executive attributes and firm outcomes such as innovation and financial performance. For example, prior research finds that gender-diverse boards tend to exhibit improved internal governance (e.g., better oversight and attendance) although the average impact on firm profitability or market value is mixed (Adams and Ferreira,, 2009). A comprehensive meta-analysis by Post and Byron, (2015) reports that female board representation is positively related to accounting returns, especially in societies with greater gender parity, but the correlation with market-based performance metrics is weaker. More recent work has begun to address endogeneity in this relationship: Safiullah et al., (2022), analyzing Spain’s Gender Equality Act, use GMM techniques and find that while gender-diverse boards outperform on accounting measures, they can underperform on market valuation measures, suggesting investors may respond differently than internal metrics. Similarly, a study of Russian firms by Safiullah et al., (2022) finds that gender-diverse boards are associated with higher profitability and market value, with the benefits particularly pronounced during economic downturns. Beyond gender, other aspects of diversity have been linked to innovation outcomes: Østergaard et al., (2011) show that employee gender and educational diversity positively predict firm innovation, and in a study of London firms, Nathan and Lee, (2013) find cultural diversity in management boosts product innovation and entrepreneurship. Field experiments also echo these benefits—Hoogendoorn et al., (2013) conducted a randomized experiment with startup teams and found that gender-balanced teams outperformed male-dominated teams in terms of sales and profits.
@@
An intriguing hypothesis within this literature is the existence of non-linear effects or “critical mass” thresholds in the diversity–performance relationship. Sociologist Rosabeth Kanter’s classic work on tokenism (Kanter,, 1977, 1987) theorized that women in extreme minority (a “token” few) face marginalization, whereas once a minority group reaches a substantial share of the team, dynamics shift and their influence grows disproportionately. Kanter’s typology categorizes group gender composition into skewed (up to $\sim$15% women), tilted ($\sim$20–35% women), and balanced ($\sim$40–50% women) categories, proposing that performance benefits might emerge when moving from skewed to tilted or balanced distributions. Subsequent studies have sought empirical evidence of such tipping points. For instance, Torchia et al., (2011) find that having at least three women directors (roughly a critical mass on many boards) is associated with a jump in innovation outputs, consistent with moving beyond token representation. Ali et al., (2011) report an inverted U-shaped relationship between female representation and firm performance in certain contexts, suggesting that the strongest returns may occur at intermediate diversity levels before tapering off. These studies, while suggestive, largely report correlations or rely on linear/quadratic models that may not capture the true causal threshold. Our study contributes to this literature by using a robust, partially identified approach to formally test for causal tipping points. By refraining from imposing a specific functional form, we let the data reveal whether and where increasing diversity has a statistically reliable positive effect on firm value.
@@
Our empirical analysis uses a panel of 945 publicly listed firms observed quarterly from 2015 Q2 to 2022 Q1. We focus on Tobin’s $Q$ (the ratio of market value to the replacement cost of assets) as the outcome of interest, which is a standard proxy for a firm’s growth opportunities and innovative performance. Originally introduced by Tobin, (1969) and later expounded in Tobin’s subsequent work (Tobin,, 1978), the $Q$-ratio captures market expectations of future returns—a value above 1 indicates that the firm’s market valuation exceeds book value, signaling strong investment incentives (Brainard and Tobin,, 1968; Tobin and Brainard,, 1976). For each quarter, we define the “treatment” as whether the firm’s top management team or board exceeds a given diversity threshold. In separate analyses, we consider thresholds for the percentage of women in senior leadership (e.g., 30%, 40%, 50%, etc.), reflecting the critical mass levels discussed above. We then estimate the nonparametric bounds on the ATE of diversity at each threshold using our concATE procedure. This approach does not assume that firms with different diversity levels are comparable on unobservables; instead, it provides an interval estimate for the possible causal effect, given the observable data, without invoking full identification. In contrast to most prior studies that report point estimates after making identification assumptions, our results will highlight the range of plausible causal impacts of diversity on Tobin’s $Q$, emphasizing what can be learned with minimal assumptions. Our findings yield informative insights. In broad terms, the concATE analysis suggests that senior-level gender diversity has a significantly positive causal effect on Tobin’s QQ—but only after a certain threshold of representation is achieved. In innovation-driven sectors (such as technology and healthcare, where overall growth opportunities are high), we find that once female representation in leadership surpasses roughly one-third, the lower bound of the ATE becomes positive and the confidence band excludes zero. The estimated effect size grows as diversity increases, with particularly strong gains evident as teams approach gender balance (around 50% female). This provides empirical support for the notion of a “tipping point” around moderate to high diversity levels in dynamic industries. One interpretation is that innovation-oriented firms, facing fast-moving and competitive markets, have strong incentives to harness the benefits of workforce diversity. Such firms may actively invest in inclusive cultures and leadership practices that allow diverse perspectives to be heard and integrated, thereby capturing value from diversity once a basic critical mass is present. By contrast, in more traditional or cyclically oriented industries, the data suggest that a much higher critical mass—on the order of two-thirds female representation—is needed before we detect a reliably positive impact on firm value. Below that level, the confidence bands include zero, indicating we cannot rule out no effect in those sectors. This stringent “tipping point” in traditional industries may reflect a lack of inclusion; when women remain a small minority, they may not experience the psychological safety needed to freely voice their insights or challenge prevailing viewpoints. Alternatively, it is possible that the gains to diversity are lower in this context. Importantly, these conclusions are drawn with rigorous uncertainty quantification. The concATE bands allow us to assert, for example, that at 95% confidence a firm in a growth industry with a gender-balanced leadership enjoys an ATE on Tobin’s $Q$ that is positive (bounded away from zero), whereas at lower diversity levels the effect cannot be distinguished from zero. Such results illustrate how our methodological innovation can uncover nuanced causal relationships that might be obscured or misstated by conventional point estimation approaches.
@@ 2 Framework
In this section, inspired by the noted shortcomings in causal inference methodologies rigorously discussed by Angrist and Pischke, (2009), we extend the theoretical framework of Manski, (1990, 2003) to derive nonparametric bounds on the “average” diversity treatment effect.
@@ 2.1 Nonparametric Bounds
Let us denote the potential outcomes for firm $i$ in sector $j$ in quarter $t$ by $Y_{ijt}^{(0)}$ and $Y_{ijt}^{(1)}$, corresponding to the scenarios of no diversity efforts (no treatment) and with diversity efforts (treatment), respectively. Regardless of whether firm $i$ actually adopts diversity, $Y_{ijt}^{(0)}$ represents the hypothetical (counterfactual) outcome had the firm not exercised any diversity efforts, and $Y_{ijt}^{(1)}$ represents the outcome if the firm did adopt diversity. In essence, the question we seek to investigate is whether these potential outcomes differ—i.e., whether diversity efforts affect $Y_{ijt}$.
@@
where $\mathbbm{1}\{\cdot\}$ is the indicator function, $\mathcal{D}(z_{ijt})$ is a diversity signal, and $\tau$ is a threshold chosen by the investigator. Let $\mathbf{X}_{ijt}=(X_{ijt}^{1},\dots,X_{ijt}^{p})^{\top}\in\mathbb{R}^{p}$ denote a $(p\times 1)$ vector of control variables. Our goal is to learn the conditional treatment effect $Y_{ijt}^{(1)}-Y_{ijt}^{(0)}\mid\mathbf{X}_{ijt}$. Following the notation of Angrist and Pischke, (2009), the observed outcome can be written in terms of potential outcomes as
@@
$\mathbb{E}\left[Y_{ijt}\mid\mathbf{X}_{ijt},Z_{ijt}=1\right]-\mathbb{E}\left[Y_{ijt}\mid\mathbf{X}_{ijt},Z_{ijt}=0\right].$ (4)
@@
$\begin{split}\delta(\mathbf{X})=&\mathbb{E}\left[Y_{ijt}\mid\mathbf{X}_{ijt},Z_{ijt}=1\right]-\mathbb{E}\left[Y_{ijt}\mid\mathbf{X}_{ijt},Z_{ijt}=0\right]\\[4.0pt] =&\underbrace{\mathbb{E}\left[Y_{ijt}^{(1)}-Y_{ijt}^{(0)}\mid\mathbf{X}_{ijt},Z_{ijt}=1\right]}_{\rho(\mathbf{X})}\\[2.0pt] &\quad+\underbrace{\mathbb{E}\left[Y_{ijt}^{(0)}\mid\mathbf{X}_{ijt},Z_{ijt}=1\right]-\mathbb{E}\left[Y_{ijt}^{(0)}\mid\mathbf{X}_{ijt},Z_{ijt}=0\right]}_{\mathcal{B}(\mathbf{X})},\end{split}$ (5)
where $\rho(\mathbf{X})$ is the (conditional) treatment effect and $\mathcal{B}(\mathbf{X})$ is the selection bias. As a corollary, the unconditional mean‐comparison parameter $\delta$ of Angrist and Pischke, (2009) is obtained by integrating $\delta(\mathbf{X}_{ijt})$ over the distribution of $\mathbf{X}_{ijt}$. Hence
$\delta=\mathbb{E}_{\mathbf{X}\mid Z}\left[\delta(\mathbf{X})\right]=\mathbb{E}\left[Y_{ijt}\mid Z_{ijt}=1\right]-\mathbb{E}\left[Y_{ijt}\mid Z_{ijt}=0\right].$ (6)
The latter may be non‑zero because firms that adopt diversity efforts might do so precisely when they face innovation shortfalls, either to signal responsiveness to investors or to diversify their workforce in search of new ideas; in such cases $\mathcal{B}(\mathbf{X})<0$. Conversely, if a firm scales up diversity after large innovation gains, aiming to sustain that momentum, then $\mathcal{B}(\mathbf{X})>0$.
Manski, (1990, 2003) formalise the problem differently. For firms characterised by attributes $\mathbf{X}$, define the difference in expected outcomes as
@@
$\mathbb{E}\left[Y_{ijt}^{(1)}\mid\mathbf{X}_{ijt}\right]$ $=\mathbb{E}\left[Y_{ijt}^{(1)}\mid\mathbf{X}_{ijt},Z_{ijt}=1\right]\Pr\left(Z_{ijt}=1\mid\mathbf{X}_{ijt}\right)$
$\quad+\mathbb{E}\left[Y_{ijt}^{(1)}\mid\mathbf{X}_{ijt},Z_{ijt}=0\right]\Pr\left(Z_{ijt}=0\mid\mathbf{X}_{ijt}\right),$ (8)
@@
$\mathbb{E}\left[Y_{ijt}^{(1)}\mid\mathbf{X}_{ijt},Z_{ijt}=0\right]\quad\text{and}\quad\mathbb{E}\left[Y_{ijt}^{(0)}\mid\mathbf{X}_{ijt},Z_{ijt}=1\right],$ (9)
are never observed in the data. Put differently, we do not observe the innovation outcome a firm would have achieved without diversity efforts when it actually implemented them ($Z_{ijt}=1$), nor the outcome with diversity efforts when it did not implement them ($Z_{ijt}=0$).
In both scenarios, one can conduct a randomized experiment, as noted by Angrist and Pischke, (2009), which coincides with the mean-independence assumption in Manski, (2003). In that case:
$\mathbb{E}\left[Y_{ijt}^{(k)}\mid\mathbf{X}_{ijt},Z_{ijt}=1\right]=\mathbb{E}\left[Y_{ijt}^{(k)}\mid\mathbf{X}_{ijt},Z_{ijt}=0\right],\quad\text{for }k=0,1.$ (10)
@@
$\Re(\mathbf{X})=\mathbb{E}[Y_{ijt}^{(1)}\mid\mathbf{X}_{ijt},Z_{ijt}=1]-\mathbb{E}[Y_{ijt}^{(0)}\mid\mathbf{X}_{ijt},Z_{ijt}=0].$ (11)
@@
where $Q_{Y}(p)=\inf\{y:F_{Y}(y)\geq p\}$, with $F_{Y}(\cdot)$ the CDF of $Y$ and $p^{c}$ is the complement of $p$, i.e., $p^{c}=1-p$. Suppose further that we are interested in the average treatment effect rather than the effect for each unit. Using the law of iterated expectations:

### v2

# Finite-Sample Nonparametric Bounds with an Application to the Causal Effect of Workforce Gender Diversity on Firm Performance

@@ Abstract
Classical Manski bounds identify average treatment effects under minimal assumptions but, in finite samples, they rely on latent outcome expectations being bounded by the sample’s own extrema or known population bounds, an assumption often violated in firm-level data with heavy-tailed outcomes. We develop a finite-sample, concentration-driven confidence band (concATE) that replaces this requirement with a Dvoretzky-Kiefer-Wolfowitz tail bound, combines it with delta-method variance, and allocates size via a Bonferroni correction. The band extends to a group-sequential design that controls the family-wise error rate when the first “significant” diversity threshold is data-chosen. Applied to data on 901 listed firms (2015 Q2–2022 Q1), concATE shows that senior-level gender diversity has a significant positive effect on firm value (Tobin’s $Q$) only after crossing substantial representation thresholds: in Growth & Innovation sectors the effect becomes statistically significant at the 5% level once women hold roughly 55% of senior leadership roles, whereas in Defensive sectors a significant impact appears only once female leadership reaches about 60%.
@@ 1 Introduction
Estimating causal effects in settings with partially unobserved counterfactuals is a fundamental challenge in econometrics. Whenever only one of two potential outcomes is observed for each unit, the average treatment effect (ATE) cannot be point-identified without additional assumptions. Recognizing this, a stream of research following Manski’s seminal work has developed nonparametric bounds for causal effects under minimal assumptions (manski1990nonparametric; manski2003partial). The classical Manski bounds make virtually no assumptions beyond knowing the treatment status and an outcome bound, instead asking: how large or small could the true ATE be, given the data we actually observe? While this worst-case approach guarantees partial identification under arbitrary heterogeneity and certain forms of selection on unobservables, it comes at the cost of wide intervals. In policy settings where the cost of acting on a wrong sign is high, such honesty can be preferable to a potentially misleading precise estimate. However, Manski’s bounds have a critical limitation in finite samples: they implicitly assume the unseen counterfactual outcomes lie within known extremes (e.g. the sample minima and maxima or exogenously given bounds). In practice, especially with heavy-tailed outcomes like firm performance, this assumption is often violated. If the true outcome distribution extends beyond the observed range, the traditional bounds can severely undercover the true effect or even become uninformative (infinite) when no credible global bound is available. In short, classical bounds that are valid asymptotically may fail to cover the true ATE in finite samples when outcomes are unbounded. This exposes a methodological gap: how can we conduct inference on partially identified effects without assuming away heavy-tail risks or sacrificing finite-sample validity?
@@
We address this gap by proposing a finite-sample, concentration-driven confidence band for the ATE (henceforth concATE). The concATE methodology replaces Manski’s reliance on known outcome bounds with a probabilistic concentration bound that accounts for sampling uncertainty in extreme values. Specifically, we exploit the Dvoretzky-Kiefer-Wolfowitz (DKW) inequality (dvoretzky1956asymptotic), which provides a tight, distribution-free bound on the maximum discrepancy between the empirical distribution and the true cumulative distribution. By using the DKW inequality, we can guarantee with high finite-sample probability that the empirical range (or other tail statistics) bounds the latent outcome distribution. In essence, instead of assuming the sample extremes equal the population extremes, we allow a margin such that each unobserved tail probability is covered with finite-sample confidence. We then incorporate this “DKW padding” into the estimation of Manski’s upper and lower bound components. To account for sampling variability in the observable parts (like the treated and untreated outcome means), we employ standard delta-method approximations. Finally, we combine these elements using Bonferroni’s inequality to construct a simultaneous confidence band for the ATE bounds. This concATE band controls the family-wise error rate for the entire interval estimate, ensuring that with (for example) 95% confidence the true ATE lies within the band.
@@
Notably, the concATE procedure remains valid under quite general conditions: we require no parametric outcome distribution, only mild tail assumptions and allow for either independent or weakly dependent observations (such as time-series panels) with appropriate mixing conditions. The resulting inference is robust in finite samples, avoiding the need for large-sample approximations or unknown nuisance constants that plague fully nonparametric approaches. By construction, concATE eliminates the strong functional form and ignorability assumptions that conventional regression-based analyses demand, delivering credible inference even when treatment assignment may be endogenous or outcomes are highly non-normal. In contrast to an OLS or panel regression that produces a single point estimate under strict assumptions (angrist2009mostly), our approach yields a range of plausible effects consistent with the data and lets the data speak when identification is weak. In summary, concATE provides a new tool for causal inference under partial observability, offering the transparency of Manski’s bounds with greater practical applicability in finite samples.
@@
To demonstrate the utility of our approach, we apply it to the question: Does greater senior-level gender workforce diversity causally improve firm performance? This question has taken on renewed importance as many firms have invested heavily in Diversity, Equity, and Inclusion (DEI) initiatives, yet establishing causality is difficult due to non-random adoption of diversity practices. A rich literature in management and economics has examined links between top management team composition and organizational outcomes. The foundational “upper echelons” theory of hambrick1984upper posits that a firm’s strategies and performance reflect the backgrounds of its senior executives. Consistent with this view, numerous studies document associations between executive attributes and firm outcomes such as innovation and financial performance. For example, prior research finds that gender-diverse boards tend to exhibit improved internal governance (e.g., better oversight and attendance) although the average impact on firm profitability or market value is mixed (adams2009women). A comprehensive meta-analysis by post2015women reports that female board representation is positively related to accounting returns, especially in societies with greater gender parity, but the correlation with market-based performance metrics is weaker. More recent work has begun to address endogeneity in this relationship: safiullah2022gender, analyzing Spain’s Gender Equality Act, use GMM techniques and find that while gender-diverse boards outperform on accounting measures, they can underperform on market valuation measures, suggesting investors may respond differently than internal metrics. Similarly, a study of Russian firms by safiullah2022gender finds that gender-diverse boards are associated with higher profitability and market value, with the benefits particularly pronounced during economic downturns. Beyond gender, other aspects of diversity have been linked to innovation outcomes: ostergaard2011does show that employee gender and educational diversity positively predict firm innovation, and in a study of London firms, nathan2013cultural find cultural diversity in management boosts product innovation and entrepreneurship. Field experiments also echo these benefits. hoogendoorn2013impact conducted a randomized experiment with startup teams and found that gender-balanced teams outperformed male-dominated teams in terms of sales and profits.
@@
An intriguing hypothesis within this literature is the existence of non-linear effects or “critical mass” thresholds in the diversity-performance relationship. Sociologist Rosabeth Kanter’s classic work on tokenism (kanter1977some; kanter1987men) theorized that women in extreme minority (a “token” few) face marginalization, whereas once a minority group reaches a substantial share of the team, dynamics shift and their influence grows disproportionately. Kanter’s typology categorizes group gender composition into skewed (up to $\sim$15% women), tilted ($\sim$20–35% women), and balanced ($\sim$40–50% women) categories, proposing that performance benefits might emerge when moving from skewed to tilted or balanced distributions. Subsequent studies have sought empirical evidence of such tipping points. For instance, torchia2011women find that having at least three women directors (roughly a critical mass on many boards) is associated with a jump in innovation outputs, consistent with moving beyond token representation. ali2011gender report an inverted U-shaped relationship between female representation and firm performance in certain contexts, suggesting that the strongest returns may occur at intermediate diversity levels before tapering off. These studies, while suggestive, largely report correlations or rely on linear/quadratic models that may not capture the true causal threshold. Our study contributes to this literature by using a robust, partially identified approach to formally test for causal tipping points. By refraining from imposing a specific functional form, we let the data reveal whether and where increasing diversity has a statistically reliable positive effect on firm value.
@@
Our empirical analysis uses a panel of 901 publicly listed firms observed quarterly from 2015 Q2 to 2022 Q1. We focus on Tobin’s $Q$ (the ratio of market value to the replacement cost of assets) as the outcome of interest, which is a standard proxy for a firm’s growth opportunities and innovative performance. Originally introduced by tobin1969general and later expounded in Tobin’s subsequent work (tobin1978monetary), the $Q$-ratio captures market expectations of future returns. A value above 1 indicates that the firm’s market valuation exceeds book value, signalling strong investment incentives (brainard1968pitfalls; tobin1976asset). For each quarter, we define the “treatment” as whether the firm’s top management team or board exceeds a given diversity threshold. In separate analyses, we consider thresholds for the percentage of women in senior leadership (e.g., 30%, 40%, 50%, etc.), reflecting the critical mass levels discussed above. We then estimate the nonparametric bounds on the ATE of diversity at each threshold using our concATE procedure. This approach does not assume that firms with different diversity levels are comparable on unobservables; instead, it provides an interval estimate for the possible causal effect, given the observable data, without invoking full identification. In contrast to most prior studies that report point estimates after making identification assumptions, our results will highlight the range of plausible causal impacts of diversity on Tobin’s $Q$, emphasizing what can be learned with minimal assumptions.
Our findings yield informative insights. In broad terms, the concATE analysis suggests that senior-level gender diversity has a significantly positive causal effect on Tobin’s $Q$, but only after a certain threshold of representation is achieved. In innovation-driven sectors (such as technology and healthcare, where overall growth opportunities are high), we find that once female representation in leadership surpasses roughly 55%, the lower bound of the ATE becomes positive and the confidence band excludes zero. The estimated effect size grows as diversity increases, with particularly strong gains evident as teams approach and exceed gender balance (around 50% female). This provides causal empirical support for the notion of a “tipping point” around high diversity levels in dynamic industries. One interpretation is that innovation-oriented firms, facing fast-moving and competitive markets, have strong incentives to harness the benefits of workforce diversity. Such firms may actively invest in inclusive cultures and leadership practices that allow diverse perspectives to be heard and integrated, thereby capturing value from diversity once a basic critical mass is present.
By contrast, in more traditional or cyclically oriented industries, the data suggest that a higher critical mass (on the order of 60% female representation) is needed before we detect a reliably positive impact on firm value. Below that level, the confidence bands include zero, indicating we cannot rule out no effect in those sectors (but we can rule out negative effects). This stringent “tipping point” in traditional industries may reflect a lack of inclusion—when women remain a small minority, they may not experience the psychological safety needed to freely voice their insights or challenge prevailing viewpoints. Indeed, diversity alone (without an inclusive environment) can lead to friction or marginalization of minority members, whereas inclusion actively involves and values those members. This aligns with prior research suggesting that diversity must be complemented by inclusion to unlock its benefits: simply adding a few token individuals from underrepresented groups often fails to improve outcomes unless the organizational climate empowers those individuals to participate fully (roberson2006disentangling; nishii2013benefits; josten2025makes). Alternatively, it is possible that the intrinsic gains to diversity are lower in these contexts. Importantly, these conclusions are drawn with rigorous uncertainty quantification. The concATE bands allow us to assert, for example, that at 95% confidence a firm in a growth industry with a gender-balanced leadership enjoys an ATE on Tobin’s $Q$ that is positive (bounded away from zero), whereas at lower diversity levels the effect cannot be distinguished from zero. Such results illustrate how our methodological innovation can uncover nuanced causal relationships that might be obscured or misstated by conventional point estimation approaches.
@@ 2 Framework
In this section, inspired by the noted shortcomings in causal inference methodologies rigorously discussed by angrist2009mostly, we extend the theoretical framework of manski1990nonparametric; manski2003partial to derive nonparametric bounds on the “average” diversity treatment effect.
@@ 2.1 Nonparametric Bounds
Let us denote the potential outcomes for firm $i$ in sector $j$ in quarter $t$ by $Y_{ijt}^{(0)}$ and $Y_{ijt}^{(1)}$, corresponding to the scenarios of no diversity efforts (no treatment) and with diversity efforts (treatment), respectively. Regardless of whether firm $i$ actually adopts diversity, $Y_{ijt}^{(0)}$ represents the hypothetical (counterfactual) outcome had the firm not exercised any diversity efforts, and $Y_{ijt}^{(1)}$ represents the outcome if the firm did adopt diversity. In essence, the question we seek to investigate is whether these potential outcomes differ (i.e., whether diversity efforts affect $Y_{ijt}$).
@@
where $\mathbbm{1}\{\cdot\}$ is the indicator function, $\mathcal{D}$ is a diversity signal, and $\tau$ is a threshold chosen by the investigator. Let $\mathbf{X}_{ijt}=(X_{ijt}^{1},\dots,X_{ijt}^{p})^{\top}\in\mathbb{R}^{p}$ denote a $(p\times 1)$ vector of control variables. Our goal is to learn the conditional treatment effect $Y_{ijt}^{(1)}-Y_{ijt}^{(0)}\mid\mathbf{X}_{ijt}$. Following the notation of angrist2009mostly, the observed outcome can be written in terms of potential outcomes as
@@
$\delta(\mathbf{X}):=\mathbb{E}\left[Y_{ijt}\mid\mathbf{X}_{ijt},Z_{ijt}(\tau)=1\right]-\mathbb{E}\left[Y_{ijt}\mid\mathbf{X}_{ijt},Z_{ijt}(\tau)=0\right].$ (4)
@@
$\begin{split}\delta(\mathbf{X}):=&\underbrace{\mathbb{E}\left[Y_{ijt}^{(1)}-Y_{ijt}^{(0)}\mid\mathbf{X}_{ijt},Z_{ijt}(\tau)=1\right]}_{\rho(\mathbf{X})}\\[2.0pt] &\quad+\underbrace{\mathbb{E}\left[Y_{ijt}^{(0)}\mid\mathbf{X}_{ijt},Z_{ijt}(\tau)=1\right]-\mathbb{E}\left[Y_{ijt}^{(0)}\mid\mathbf{X}_{ijt},Z_{ijt}(\tau)=0\right]}_{\mathcal{B}(\mathbf{X})},\end{split}$ (5)
where $\rho(\mathbf{X})$ is the (conditional) treatment effect and $\mathcal{B}(\mathbf{X})$ is the selection bias. As a corollary, the unconditional mean‐comparison parameter $\delta$ of angrist2009mostly is obtained by integrating $\delta(\mathbf{X})$ over the (marginal) distribution of $\mathbf{X}\mid Z(\tau)$, i.e.,
$\begin{split}\delta=\int_{\mathbb{R}^{p}}&\mathbb{E}\big[Y\mid\mathbf{X},Z(\tau)=1\big]f_{\mathbf{X}\mid Z=1}(\mathbf{x})\,\mathrm{d}\mathbf{x}\\ &\quad-\int_{\mathbb{R}^{p}}\mathbb{E}\big[Y\mid\mathbf{X},Z(\tau)=0\big]f_{\mathbf{X}\mid Z=0}(\mathbf{x})\mathrm{d}\mathbf{x},\end{split}$
hence
$\delta=\mathbb{E}\left[Y_{ijt}\mid Z_{ijt}(\tau)=1\right]-\mathbb{E}\left[Y_{ijt}\mid Z_{ijt}(\tau)=0\right].$ (6)
The latter may be non-zero because firms that adopt diversity efforts might do so precisely when they face innovation shortfalls, either to signal responsiveness to investors or to diversify their workforce in search of new ideas; in such cases $\mathcal{B}(\mathbf{X})<0$. Conversely, if a firm scales up diversity after large innovation gains, aiming to sustain that momentum, then $\mathcal{B}(\mathbf{X})>0$.
manski1990nonparametric; manski2003partial formalise the problem differently. For firms characterised by attributes $\mathbf{X}$, define the difference in expected outcomes as
@@
$\mathbb{E}\left[Y_{ijt}^{(1)}\mid\mathbf{X}_{ijt}\right]$ $=\mathbb{E}\left[Y_{ijt}^{(1)}\mid\mathbf{X}_{ijt},Z_{ijt}(\tau)=1\right]\Pr\left(Z_{ijt}=1\mid\mathbf{X}_{ijt}\right)$
$\quad+\mathbb{E}\left[Y_{ijt}^{(1)}\mid\mathbf{X}_{ijt},Z_{ijt}(\tau)=0\right]\Pr\left(Z_{ijt}=0\mid\mathbf{X}_{ijt}\right),$ (8)
@@
$\mathbb{E}\left[Y_{ijt}^{(1)}\mid\mathbf{X}_{ijt},Z_{ijt}(\tau)=0\right]\quad\text{and}\quad\mathbb{E}\left[Y_{ijt}^{(0)}\mid\mathbf{X}_{ijt},Z_{ijt}(\tau)=1\right],$ (9)
are never observed in the data. Put differently, we do not observe the innovation outcome a firm would have achieved without diversity efforts when it actually implemented them ($Z_{ijt}(\tau)=1$), nor the outcome with diversity efforts when it did not implement them ($Z_{ijt}(\tau)=0$).
In both scenarios, one can conduct a randomized experiment, as noted by angrist2009mostly, which coincides with the mean-independence assumption in manski2003partial. In that case:
$\mathbb{E}\left[Y_{ijt}^{(k)}\mid\mathbf{X}_{ijt},Z_{ijt}(\tau)=1\right]=\mathbb{E}\left[Y_{ijt}^{(k)}\mid\mathbf{X}_{ijt},Z_{ijt}(\tau)=0\right],\quad\text{for }k=0,1.$ (10)
@@
$\Re(\mathbf{X})=\mathbb{E}[Y_{ijt}^{(1)}\mid\mathbf{X}_{ijt},Z_{ijt}(\tau)=1]-\mathbb{E}[Y_{ijt}^{(0)}\mid\mathbf{X}_{ijt},Z_{ijt}(\tau)=0].$ (11)
@@
where $Q_{Y}(p)=\inf\{y:F_{Y}(y)\geq p\}$, with $F_{Y}(\cdot)$ the CDF of $Y$ and $p^{c}$ is the complement of $p$, i.e., $p^{c}=1-p$. Suppose further that we are interested in the unconditional average treatment effect rather than the effect for each unit. Using the law of iterated expectations: